OpenStudy (anonymous):

* * * {LAST QUESTION; MEDAL} * * * A study of one thousand teens found that the number of hours they spend on social networking sites each week is normally distributed with a mean of 15 hours. The population standard deviation is 2 hours. What is the 95% confidence interval for the mean? 14.88−15 hours 14.88−15.12 hours 15−15.12 hours 14.76−14.24 hours

OpenStudy (anonymous):

I cant help but here is someone who might be able to: @hartnn

OpenStudy (tkhunny):

Do we have an Empirical Rule? You may wish to try two standard deviations on either side of the mean.

OpenStudy (anonymous):

Yes, 68% at 1 and -1 95% at 2 and -2 99.7% at 3 and -3 But for 95% from the standard deviation, i got 11 and 19, which isn't one of the options..?

OpenStudy (tkhunny):

Have you considered that it wants the interval of the MEAN? You may wish to use \(SD_{Mean} = \dfrac{2}{\sqrt{1000}}\)

OpenStudy (anonymous):

15 = 2/31.6227 What do I do with this?

OpenStudy (tkhunny):

How did you get 11-19? It's the same thing, just with the right Standard Deviation.

OpenStudy (anonymous):

What is the right standard deviation? Sorry, I don't understand.

OpenStudy (tkhunny):

Did you miss \(SD_{Mean}\) that I supplied, above?

OpenStudy (anonymous):

No, but I just don't know what to do with that. It would be 2^15?

OpenStudy (tkhunny):

Reset!!!! How did you find 11-19? EXACTLY what did you do to get that?

OpenStudy (anonymous):

Add and subtract the SD from both sides of the mean.

OpenStudy (anonymous):

What was I supposed to do?

OpenStudy (tkhunny):

Why not just do that again? You used \(SD_{Population} = 2\) Do it again with \(SD_{Mean} = \dfrac{2}{\sqrt{1000}}\)

OpenStudy (anonymous):

oh, like add/subtract 30 from both sides?

OpenStudy (tkhunny):

No, the Standard Deviation of the Mean is \(\dfrac{2}{\sqrt{1000}} = 0.0632455532\) Where are you getting 30?

OpenStudy (anonymous):

I multiplied 2 and 15. So on the right side, I add 0.063 and subtract that number on the left? From the mean?

OpenStudy (tkhunny):

Why are you doing ANYTHING with 2 after calculating the Standard Deviation of the Mean? The center of your confidence interval is 15. Never change that. The Standard deviation is 0.0632455532. Go!

OpenStudy (anonymous):

Ok this is just confusing, sorry.....so, 15.0632, 15.1264, 15.1896 on the left side 14.9368, 14.8736, 14.8104 on the right?

OpenStudy (anonymous):

Oh I see now. Thx for the help.